OpenStudy (anonymous):

last question! a baseball if thrown almost straight up with a speed of 20m/s. How long is it in the air?

OpenStudy (anonymous):

20m

OpenStudy (anonymous):

4seconds. vf=vo-gt 20=-20-10t Therefore t=4s

OpenStudy (anonymous):

thank you!